# Hierarchical Correlation Clustering and Tree Preserving Embedding

### v2

@@ 3 Hierarchical Correlation Clustering
We emphasize that HCC is consistent with the generic agglomerative clustering framework applied with, for example, single linkage, average linkage, complete linkage and other criteria. The only difference is the definition of the inter-cluster dissimilarity function where with HCC we use $dis^{CC}(.,.)$ defined in Eq. 19 (inspired from the cost function of flat correlation clustering). Other than this, the algorithmic procedure is consistent.
@@
Algorithm 1 describes hierarchical correlation clustering (HCC) in detail. The algorithm at the beginning assumes $n$ singleton clusters, one for each object. For each cluster, it obtains the nearest cluster and the respective similarity. The algorithm then iteratively performs the following steps. i) Finds the two nearest clusters according to the inter-cluster (dis)similarity function defined in Eq. 19. ii) Merges the respective clusters to build a new cluster at a higher level. iii) Updates the inter-cluster similarity matrix $\mathbf{S}$, the nearest neighbor vector $nn\_ind$ and the respective similarities $nn\_sim$.  [^3] In our implementation, we use a data structure similar to the linkage matrix used by scipy package in Python to encode the dendrogram and store the intermediate clusters.
@@
4 Feature Extraction from HCC
As mentioned, HCC represents the relations between objects according to the way they join to form the hierarchy. In this section, we use this intuition and develop a data representation consistent with HCC. In particular, we adapt the methods in [65, 66, 18] to our setting. For this purpose, we first introduce distance functions over a dendrogram, and then, investigate the embedding of such a distance function. This procedure leads to obtaining a set of features from HCC for each object which then can be used in the downstream task.
4.1 Distance functions over a HCC dendrogram
Given dendrogram $T$, each cluster $\mathbf{v}\in T$ represents the root of a dendrogram $T^{\prime}\in\mathcal{T}_{T}$. $T^{\prime}$ takes the properties of its root cluster, i.e., $level(T^{\prime})=\max_{\mathbf{u}\in T^{\prime}}level(\mathbf{u})=level(\mathbf{v})$ and $linkage(T^{\prime})=\max_{\mathbf{u}\in T^{\prime}}linkage(\mathbf{u})=linkage(\mathbf{v})$, since the root cluster has the maximum linkage and level among the clusters in $T^{\prime}$. Hence, in this way, we define functions such as $level()$ and $linkage()$ for the dendrograms as well.
The $linkage()$ function seems to be a natural choice for defining a distance function over a dendrogram. Specifically, one can define the dendrogram-based distance function $\mathbf{X}_{ij}$ over dendrogram $T$ between $i,j\in\mathbf{O}$ as
This choice corresponds to the linkage of the smallest cluster that includes both $i$ and $j$. This in particular makes sense for the single linkage dendrogram, and it would be consistent with the tree-preserving embedding in [65, 66]. If the original dissimilarity matrix $\mathbf{D}$ contains negative values, then using Proposition 1, one can sufficiently shift the pairwise dissimilarities to make all of them non-negative, without changing the structure of the dendrogram and the order of the clusters. Therefore, the conditions for a valid distance function (i.e., non-negativity) are still fulfilled.
However, for the HCC dendrogram, the linkage function might not satisfy the conditions for a distance function. For example, consider a set of $n$ objects where all the pairwise similarities are $+1$, i.e., the dissimilarities are thus $-1$. Then, the linkage function will always return negative values which would violate the non-negativity condition of a valid distance function. On the other hand, the HCC linkage $dis^{CC}()$ is not shift-invariant (similar to the standard flat correlation clustering [33]) and we cannot use the shift trick in Proposition 1. Let $\mathbf{D}^{\alpha}$ shows the shifted pairwise dissimilarities, i.e., $\mathbf{D}^{\alpha}_{i,j}=\mathbf{D}_{i,j}+\alpha$. Then, $dis^{CC}(\mathbf{u},\mathbf{v})$ between two clusters $\mathbf{u}$ and $\mathbf{v}$ based on $\mathbf{D}^{\alpha}$ is given by
With $\alpha>0$, this shift would induce a bias for the HCC linkage to choose imbalanced clusters. In other words, $dis^{CC}()$ is not shift invariant and we cannot shift the pairwise dissimilarities in $\mathbf{D}$ to make them nonnegative.
4.2 Embedding the HCC-based distances
The distance function specified in Eq. 22 yields an $n\times n$ matrix representing pairwise dendrogram-based distances among objects. However, numerous machine learning algorithms operate on vector-based object representations rather than pairwise distances. Models like Gaussian Mixture Models (GMMs), which involve mixture density estimation, fall into this category.
Vectors offer a fundamental data representation, establishing a one-to-one mapping between objects and measurements. This property enables the application of a wide range of machine learning methods. Additionally, working with vector-based data simplifies feature selection. Hence, it is desired to compute an embedding of the objects into a new space, so that their pairwise squared Euclidean distances in the new space match their pairwise distances obtained from the dendrogram.
We note that the matrix of pairwise distances $\mathbf{X}$ obtained via Eq. 22 induces an ultrametric [47]. The primary distinction between a metric and an ultrametric is that the addition operation in the triangle inequality for a metric is replaced by a maximum operation, i.e., with ultrametric we have
It is evident that when $\mathbf{X}_{ij}\leq\mathbf{X}_{ik}$, the inequality in Eq. 23 is satisfied. Conversely, if $\mathbf{X}_{ij}>\mathbf{X}_{ik}$, it implies that objects $i$ and $k$ are included in the same cluster (shown by $c_{i,k}$) before $i$ and $j$ join (to form cluster $c_{i,j}$). The bottom-up hierarchical clustering process then continues until $c_{i,j,k}$ is formed, encompassing all three objects $i,j,k$. Notice that $i$ and $k$ join $j$ simultaneously via $c_{i,k}$. In this case, according to Eqs. 22 and 23, and the relationships illustrated in Figure 1, we conclude
Figure 1: Illustration of ultrametric property of $\mathbf{X}$.

### v4

@@ 3 Hierarchical Correlation Clustering
We emphasize that HCC is consistent with the generic agglomerative clustering framework applied with, for example, single linkage, average linkage, complete linkage and other criteria. The only difference is the definition of the inter-cluster dissimilarity function where with HCC we use $dis^{CC}(.,.)$ defined in Eq. 5 (inspired from the cost function of flat correlation clustering). Other than this, the algorithmic procedure is consistent. Algorithm 1 in Appendix A describes the pseudocode of the HCC algorithm.
4 Feature Extraction from HCC
As mentioned, HCC represents the relations between objects according to the way they join to form the hierarchy. In this section, we use this intuition and develop a data representation consistent with HCC. For this purpose, we adapt the methods in [69, 70] and in particular [21] to our setting. Hereby, we first introduce distance functions over HCC, and then, investigate the embedding of such a distance function. This procedure leads to obtaining a set of features from HCC for each object which then can be used in the downstream task.
4.1 Distance functions over HCC
Given dendrogram $T$, each cluster $\mathbf{v}\in T$ represents the root of a dendrogram $T^{\prime}\in\mathcal{T}_{T}$. $T^{\prime}$ admits the properties of its root cluster, i.e., $level(T^{\prime})=\max_{\mathbf{u}\in T^{\prime}}level(\mathbf{u})=level(\mathbf{v})$ and $linkage(T^{\prime})=\max_{\mathbf{u}\in T^{\prime}}linkage(\mathbf{u})=linkage(\mathbf{v})$, since the root cluster has the maximum linkage and level among the clusters in $T^{\prime}$. Hence, in this way, we define functions such as $level()$ and $linkage()$ for the dendrograms as well.
The $linkage()$ function may seem to be a natural choice for defining a distance function over a HCC dendrogram. Specifically, one can define the dendrogram-based distance function $\mathbf{X}_{ij}$ over dendrogram $T$ between $i,j\in\mathbf{O}$ as
This choice corresponds to the linkage of the smallest cluster that includes both $i$ and $j$. This in particular makes sense for the single linkage dendrogram, and it would be consistent with the tree-preserving embedding in [69, 70]. If the original dissimilarity matrix $\mathbf{D}$ contains negative values, then using Proposition 1, one can sufficiently shift the pairwise dissimilarities to make all of them non-negative, without changing the structure of the dendrogram and the order of the clusters. Therefore, the conditions for a valid distance function including non-negativity still hold.
However, for the HCC dendrogram, the linkage function might not fulfill the conditions for a distance function. For example, consider a set of $n$ objects where all the pairwise similarities are $+1$, i.e., the dissimilarities are thus $-1$. Then, the linkage function will always return negative values which would violate the non-negativity condition of a valid distance function. On the other hand, the HCC linkage $dis^{CC}(.,.)$ is not shift-invariant (similar to the standard flat correlation clustering [18]) and we cannot use the shift trick in Proposition 1. Let $\mathbf{D}^{\alpha}$ shows the shifted pairwise dissimilarities, i.e., $\mathbf{D}^{\alpha}_{i,j}=\mathbf{D}_{ij}+\alpha$. Then, $dis^{CC}(\mathbf{u},\mathbf{v})$ between two clusters $\mathbf{u}$ and $\mathbf{v}$ based on $\mathbf{D}^{\alpha}$ is given by
With $\alpha>0$, this shift would induce a bias for the HCC linkage to choose imbalanced clusters. In other words, $dis^{CC}(.,.)$ is not shift-invariant and we cannot shift the pairwise dissimilarities in $\mathbf{D}$ to make them nonnegative.
4.2 Embedding the HCC-based distances
After applying the distance function in Eq. 8, we obtain an $n\times n$ matrix representing pairwise HCC-based distances among objects. It is usually preferred to obtain vector-based representations for objects rather than pairwise distances. Models like Gaussian Mixture Models (GMMs) which involve mixture density estimation (see, e.g., [75]), can only be applied to vectors. Additionally, working with vector-based data simplifies feature selection. Hence, it is desired to compute an embedding of the objects into a new space, so that their pairwise squared Euclidean distances in the new space match their pairwise distances obtained from the dendrogram.
The matrix of pairwise distances $\mathbf{X}$ obtained via Eq. 8 induces an ultrametric [51, 21]. The primary distinction between a metric and an ultrametric is that the addition operation in the triangle inequality for a metric is replaced by a maximum operation, i.e., with ultrametric we have
It is evident that when $\mathbf{X}_{ij}\leq\mathbf{X}_{ik}$, the inequality in Eq. 9 is satisfied. Conversely, if $\mathbf{X}_{ij}>\mathbf{X}_{ik}$, it implies that objects $i$ and $k$ are included in the same cluster (shown by $c_{i,k}$) before $i$ and $j$ join (to form cluster $c_{i,j}$). The bottom-up hierarchical clustering process then continues until $c_{i,j,k}$ is formed, encompassing all three objects $i,j,k$. Notice that $i$ and $k$ join $j$ simultaneously via $c_{i,k}$. In this case, according to Eqs. 8 and 9, and the relationships illustrated in Figure 1, we conclude
Figure 1: Illustration of ultrametric property of $\mathbf{X}$.
  1. 1.
    We center $\mathbf{X}$ by
    $\mathbf{B}\leftarrow-\frac{1}{2}(\mathbf{I}_{n}-\frac{1}{n}\mathbf{L}_{n})\mathbf{X}(\mathbf{I}_{n}-\frac{1}{n}\mathbf{L}_{n}),$ (11)
    where $\mathbf{I}_{n}$ is an identity matrix of size $n\times n$ and $\mathbf{L}_{n}$ represents an $n\times n$ matrix filled entirely with ones. With this centering, the sum of both the rows and columns in matrix $\mathbf{B}$ becomes zero.
  2. 2.
    $\mathbf{B}=\mathbf{Y}\mathbf{Z}\mathbf{Y}^{\texttt{T}},$ (12)
  3. 3.
    $\mathbf{A}=\mathbf{Y}_{l}(\mathbf{Z}_{l})^{1/2},$ (13)
In the embedded space, the new dimensions are arranged based on their corresponding eigenvalues. One can opt to select only the most significant ones, rather than utilizing all of them. Therefore, computing such an embedding also offers the benefit of feature selection.
We also note that many clustering methods can be written in matrix factorization form via for example spectral $K$-means [28]. This induces an embedding and hence a set of relevant features. However, for general positive/negative similarity matrices no exact embedding might be feasible due to violating positive semidefiniteness. The method that we described here provides a solution to this challenge, i.e., enables us to extract features when the base pairwise similarities are positive and negative.
5 Correlation Clustering and Minimax Dissimilarities
Finally, we study minimax dissimilarities for correlation clustering, a graph-based method that corresponds to constructing a hierarchical clustering.
Given graph $\mathcal{G}(\mathbf{O},\mathbf{D})$, the minimax (MM) dissimilarity between $i$ and $j$ is defined as
$\mathbf{D}_{ij}^{MM}$ $=$ $\min_{p\in\mathcal{P}_{ij}(\mathcal{G})}\max_{1\leq l\leq|p|-1}\mathbf{D}_{p(l)p(l+1)},$ (14)
Minimax dissimilarities enable a clustering algorithm to capture the inherent patterns and manifolds in an unsupervised and nonparametric way by extracting the transitive connections [32, 16]. For example, if object $i$ is similar to object $j$, $j$ is similar to $k$, and $k$ is similar to $l$, then the minimax dissimilarity between $i$ and $l$ will be small, even though their direct dissimilarity might be large. The reason is that minimax dissimilarity finds the connectivity path $i\to j\to\cdots\to k\to l$ and connects $i$ and $l$ via this path. This property is helpful in finding elongated clusters and manifolds of arbitrary shapes in an unsupervised way.
Lemma 1
All the proofs are in Appendix B. Hence, given a dissimilarity matrix $\mathbf{D}$, one can subtract $\alpha:=min(\mathbf{D})$ from all the elements to obtain $\mathbf{D}^{\alpha}$. Then, the minimax dissimilarities can be computed from $\mathcal{G}^{\alpha}(\mathbf{O},\mathbf{D}^{\alpha})$. After computing the minimax dissimilarities from $\mathcal{G}^{\alpha}$, we may add $\alpha$ to all the pairwise minimax dissimilarities. We can obtain the minimax similarities $\mathbf{S}^{MM}_{ij}$ via $\mathbf{S}^{MM}_{ij}=-\mathbf{D}^{MM}_{ij}$, if needed.
Theorem 1
As mentioned, correlation clustering on an arbitrary similarity matrix $\mathbf{S}$ is NP-hard [8, 26]. Therefore, using minimax (dis)similarities with correlation clustering not only helps for extracting elongated complex patterns, but also yields a significant reduction in the computational complexity, i.e., from NP-hardness to a polynomial runtime.
Among the approximate algorithms proposed for correlation clustering on complete graphs with discrete weights, the method in [5] provides a 3-factor approximation. with a randomly selected unclustered object in the graph, this method greedily finds the object’s positive neighbors (those with similarity $+1$) to form a new cluster. Then, it repeats this procedure for the remaining objects. One can conclude that in the optimal solution of correlation clustering on graph $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$, only the positive neighbors of an object will be in the same cluster as the object is, i.e., interestingly the 3-factor approximation algorithm in [5] becomes exact when applied to $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$ (Theorem 2).
Theorem 2
6 Experiments
In this section, we describe our experimental results on various datasets. We compare our methods with single linkage (SL), complete linkage (CL) and average linkage (AL).333 Some criteria, e.g. centroid, median and Ward compute a representative for each cluster and then compute the inter-cluster dissimilarities by the distances between the representatives. Computing such representatives might not be feasible for possibly negative pairwise dissimilarities. Thus, we do not consider them. As mentioned, there are several improvements over these basic methods. However, such contributions are orthogonal to our contribution. Moreover, it is unclear how such improvements can be extended to the dissimilarities that can be both positive and negative. Thus, we limit our baselines to these methods which as mentioned in Proposition 1, are applicable to such data.
In Appendix C, we describe additional experimental results, in particular on datasets from other domains.
(a) [Breast Tissue]
(b) Cardio
(c) Image Segmentation
(d) ISOLET
(e) Leaf
(f) One-Hundred Plant
Figure 2: MI score of different hierarchical clustering methods applied to UCI datasets, where x-axis shows the parameter $\eta$.
{adjustwidth}
0cm0cm Breast Tissue Cardiotocography Image Segm. ISOLET Leaf One-Hun. Plant method MI Rand MI Rand MI Rand MI Rand MI Rand MI Rand SL 0.006 0.008 0.007 0.007 0.008 0.001 0.009 0.016 0.008 0.003 0.015 0.020 SL+GMM 0.093 0.077 0.120 0.135 0.239 0.250 0.192 0.174 0.155 0.161 0.083 0.077 CL 0.227 0.166 0.077 0.056 0.187 0.125 0.057 0.017 0.081 0.038 0.029 0.008 CL+GMM 0.251 0.171 0.081 0.060 0.201 0.154 0.061 0.043 0.140 0.129 0.054 0.049 AL 0.542 0.519 0.391 0.479 0.518 0.495 0.257 0.165 0.181 0.106 0.066 0.023 AL+GMM 0.550 0.513 0.422 0. 463 0.522 0.501 0.240 0.179 0.152 0.143 0.081 0.065 HCC 0.903 0.900 0.987 0.994 0.945 0.943 0.938 0.918 0.429 0.373 0.159 0.104 HCC+GMM 0.914 0.911 0.979 0.974 0.960 0.966 0.941 0.917 0.462 0.401 0.183 0.217
Table 1: Performance of different tree preserving embedding methods on UCI datasets applied with GMM.
6.1 HCC on UCI data
We first investigate the hierarchical correlation clustering on the following six UCI datasets [45]. (i) Breast Tissue: includes electrical impedance measurements of freshly excised $106$ tissue samples from the breast. The number of clusters is $6$. (ii) Cardiotocography: contains $2126$ measurements of fetal heart rate and uterine contraction features on cardiotocograms in $10$ clusters. (iii) Image Segmentation: contains $2310$ samples from images of $7$ outdoor clusters. (iv) ISOLET: $7797$ samples consisting of spoken attributes of different letters ($26$ clusters). (v) Leaf: $340$ images of leaf specimens originating from $40$ different plant species (clusters) each described by $16$ attributes. (vi) One-Hundred Plant: $1600$ samples of leafs each described by $64$ features, from in total $100$ types (clusters). The ground-truth labels are shown by $\mathbf{c}^{*}$, i.e., $\mathbf{c}^{*}_{i}$ shows the true label for object $i$. We assume an oracle reveals the pairwise similarities $\mathbf{S}$ according to the (flip) noise parameter $\eta$. If $\mathbf{c}^{*}_{i}=\mathbf{c}^{*}_{j}$ then $\mathbf{S}_{i,j}=\mathcal{U}(0,1)$ with probability $1-\eta$ and $\mathbf{S}_{i,j}=\mathcal{U}(-1,0)$ with probability $\eta$. If $\mathbf{c}^{*}_{i}\neq\mathbf{c}^{*}_{j}$ then $\mathbf{S}_{i,j}=\mathcal{U}(-1,0)$ with probability $1-\eta$ and $\mathbf{S}_{i,j}=\mathcal{U}(0,1)$ with probability $\eta$. The function $\mathcal{U}(.,.)$ returns a uniform random number within the specified range. For each $\eta$ we repeat the experiments $20$ times and report the average results. This setup provides a systematic approach to study the robustness of various methods to noise.
Figure 2 shows the results for different datasets as a function of the noise level $\eta$ w.r.t. MI. Rand scores shown in Figure 5 in Appendix C exhibit a consistent behavior. We observe that among different methods, HCC performs significantly better and produces more robust clusters w.r.t. the noise parameter $\eta$. The results on Leaf and One-Hundred Plant are worse with all the methods. The reason is that these datasets are complex, having many clusters (respectively, $40$ and $100$ clusters) and fairly a small number of objects per cluster.
6.2 Tree preserving embedding on UCI data
In the following, we investigate tree preserving embedding and feature extraction. After computing the embeddings from different hierarchical clustering methods, we apply Gaussian Mixture Model (GMM) to the extracted features and evaluate the final clustering using the ground-truth solution. This kind of embedding enables us to apply methods like GMM to positive and negative pairwise similarities, a task that was not possible before. Since the extracted features appear in the form of vectors, thus, the final clustering method is not limited to GMM, and in principle, any numerical clustering can benefit from this embedding.
Table 1 shows the results for different UCI datasets. We observe that the embeddings obtained by HCC (i.e., ‘HCC+GMM’) yield significantly better results compared to the embeddings from other methods (i.e., ‘SL+GMM’, ‘CL+GMM’ and ‘AL+GMM’). It is worth noting that the results of embeddings (e.g., ‘HCC+GMM’) typically surpass the results obtained from the hierarchical clustering alone (e.g., ‘HCC’). This observation supports the idea that employing HCC to compute an embedding (for extracting new features) for a clustering method such as GMM may be advantageous, yielding superior results compared to using HCC exclusively for clustering purposes. This verifies why tree preserving embedding can be effective in general.
6.3 Experiments on Fashion-MNIST
Fashion MNIST 1 Fashion MNIST 2
method MI Rand MI Rand
SL 0.322 0.206 0.241 0.196
SL+GMM 0.411 0.335 0.384 0.340
CL 0.403 0.293 0.546 0.379
CL+GMM 0.478 0.426 0.574 0.362
AL 0.464 0.468 0.602 0.534
AL+GMM 0.608 0.551 0.647 0.553
HCC 0.499 0.475 0.666 0.557
HCC+GMM 0.581 0.586 0.693 0.548
(a) DS1 [44, 63]
(b) DS2 [22]
(c) DS3 [1]
(d) DS4 [1]
6.4 Correlation clustering with minimax dissimilarities
Finally, we investigate the use of minimax dissimilarities with correlation clustering. As mentioned, minimax dissimilarities are especially useful for extracting elongated and complex patterns in an unsupervised way. Thereby, we apply ‘minimax + correlation clustering’ to a number of datasets with visually elongated and arbitrarily shaped clusters, and compare the results with ‘correlation clustering’ alone. The datasets are shown in Figure 3. For each dataset, we simply construct the K-nearest neighbor graph (using the ordinary Euclidean distance with a typical K such as 3). The edges in the nearest neighborhood receive a positive weight (e.g., +1), and a negative weight (e.g., -1) otherwise. We then apply either ‘minimax + correlation clustering’ or ‘correlation clustering’ to the resultant graph. We observe that for all the datasets, ‘minimax + correlation clustering’ yields the perfect clustering, i.e., MI and Rand are equal to 1. Whereas with ‘correlation clustering’ alone these scores are very low. We acknowledge the existence of other clustering methods, such as DBSCAN [30], capable of achieving perfect clustering on these datasets. However, these methods often involve crucial hyperparameters, the tuning of which can be challenging in unsupervised learning. To our knowledge, in addition to computational and theoretical benefits, ‘minimax + correlation clustering’ stands out as the sole method capable of achieving perfect results on datasets with elongated clusters of arbitrary shapes, and fulfills the following two promises: i) it eliminates the need to fix critical hyperparameters, a task often intricate in unsupervised learning, and ii) it automatically determines the correct number of clusters without requiring prior fixing.
7 Conclusion
We proposed a new hierarchical clustering method, called HCC, suitable when the (dis)similarities can be positive or negative. This method is consistent with the general algorithmic procedure for agglomerative clustering and only differs in the way the inter-cluster dissimilarity function is defined. We then considered embedding the HCC dendrograms, which provides extracting useful features to apply for example GMM for clustering positive and negative similarities. In the following, we studied the use of minimax dissimilarities with correlation clustering and showed that it yields reduction in the computational complexity, in addition to a possibility for extracting elongated manifolds. Finally, we demonstrated the effectiveness of the methods on several datasets in different settings.
Acknowledgement
The work of Morteza Haghir Chehreghani was partially supported by the Swedish Research Council VR (grant number 2023-04809) and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A HCC Algorithm
@@
Algorithm 1 describes hierarchical correlation clustering (HCC) in detail. The algorithm at the beginning assumes $n$ singleton clusters, one for each object. For each cluster, it obtains the nearest cluster and the respective similarity. The algorithm then iteratively performs the following steps. i) Finds the two nearest clusters according to the inter-cluster (dis)similarity function defined in Eq. 5. ii) Merges the respective clusters to build a new cluster at a higher level. iii) Updates the inter-cluster similarity matrix $\mathbf{S}$, the nearest neighbor vector $nn\_ind$ and the respective similarities $nn\_sim$ [^4] In our implementation, we use a data structure similar to the linkage matrix used by scipy package in Python to encode the dendrogram and store the intermediate clusters.
@@
Appendix B Proofs
B.1 Proof of Lemma 1
One can show that the pairwise minimax dissimilarities across any given graph are identical to the pairwise minimax dissimilarities present in any minimum spanning tree obtained from the same graph. The proof is similar to the maximum capacity problem [40]. Thereby, the minimax dissimilarities are obtained by
$=$ $\max_{1\leq l\leq|p_{ij}|-1}\mathbf{D}_{p(l)p(l+1)},$
where $p_{ij}$ indicates the (only) path between $i$ and $j$ on a minimum spanning tree computed on $\mathcal{G}$. To obtain the minimax dissimilarities $\mathbf{D}^{MM}_{ij}$, we can just select the maximal edge weight on the only path between $i$ and $j$ on the minimum spanning tree.
On the other hand, the single linkage method and the Kruskal’s minimum spanning tree algorithm are equivalent [35]. Thus, the dendrogram $T$ obtained via single linkage sufficiently contains the pairwise minimax dissimilarities. Now, we elaborate that the minimax dissimilarities in Eq. 15 equal the dissimilarities defined in Eq. 6, i.e., $\mathbf{X}_{ij}$ is the largest edge weight on the path between $i$ and $j$ in the hierarchy.
Given $i,j$, let
B.2 Proof of Theorem 2